OpenStudy (anonymous):

3x^2+23x-8 please help me, I need this done in less than five minutes, and im lost

OpenStudy (anonymous):

It wants me to fiind x intercepts

jhonyy9 (jhonyy9):

3x^2 +15x +8x -8

OpenStudy (anonymous):

set y=0 to find x-intercepts

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

3x^2+23x-8=0

OpenStudy (anonymous):

-8, 1/3? is this the answers

OpenStudy (anonymous):

*are these

OpenStudy (anonymous):

You can use the ac method to group here. 3x^2+24x-x-8=0 3x(x+8)-(x+8)=0 (3x-1)(x+8)=0

OpenStudy (anonymous):

Yes, correct answers.

OpenStudy (anonymous):

:D